# Hierarchical filtrations of torsion-free sheaves and birational geometry

### v1

@@ 4. Hierarchical depth and determinant bounds
where each quotient is torsion supported on an effective divisor. Since $\mathcal{E}^{\prime}\subset\mathcal{E}$ is saturated of the same rank, each inclusion $\mathcal{E}^{\prime}_{i}\subset\mathcal{E}$ remains saturated in $\mathcal{E}$ and preserves the torsion quotients. Hence the same chain is a valid hierarchical filtration of $\mathcal{E}$. This proves $h_{\Lambda_{0}}(\mathcal{E}^{\prime})\leq h_{\Lambda_{0}}(\mathcal{E})$.
(B) Let $X=C$ be a smooth projective curve and normalize by $\Lambda_{0}=\mathcal{O}_{C}$. Monotonicity of hierarchical depth need not hold when passing to a proper subbundle of strictly smaller rank: one can have a subbundle $\mathcal{E}^{\prime}\subset\mathcal{E}$ with $\mathrm{rank}(\mathcal{E}^{\prime})<\mathrm{rank}(\mathcal{E})$ yet
$h(\mathcal{E}^{\prime})>h(\mathcal{E}).$
A minimal explicit example on $\mathbb{P}^{1}$ is as follows.
$\mathcal{E}^{\prime}=\mathcal{O}_{\mathbb{P}^{1}}(1),\qquad\mathcal{E}=\mathcal{O}_{\mathbb{P}^{1}}(1)\oplus\mathcal{O}_{\mathbb{P}^{1}}(-1),\qquad\mathcal{E}^{\prime}\hookrightarrow\mathcal{E}\text{ the summand inclusion.}$
Then
$\det(\mathcal{E})=\mathcal{O}_{\mathbb{P}^{1}}(1)\otimes\mathcal{O}_{\mathbb{P}^{1}}(-1)\cong\mathcal{O}_{\mathbb{P}^{1}},$
so by Proposition 4.1,
$h(\mathcal{E})=\deg\det(\mathcal{E})=0,$
whereas
$h(\mathcal{E}^{\prime})=\deg\mathcal{O}_{\mathbb{P}^{1}}(1)=1.$
Therefore $h(\mathcal{E}^{\prime})=1>0=h(\mathcal{E})$, showing that hierarchical depth is not monotone under subbundles of smaller rank.
(C) Let
$0\to\mathcal{E}^{\prime}\xrightarrow{\iota}\mathcal{E}\xrightarrow{\pi}\mathcal{E}^{\prime\prime}\to 0$
be exact. By assumption (i) there exists a normalized hierarchical filtration of $\mathcal{E}^{\prime}$,
$\mathcal{E}^{\prime}_{0}\subset\mathcal{E}^{\prime}_{1}\subset\cdots\subset\mathcal{E}^{\prime}_{a}=\mathcal{E}^{\prime},$
with $\det(\mathcal{E}^{\prime}_{0})\simeq\Lambda_{0}$ and determinant increments $\det(\mathcal{E}^{\prime}_{i})\simeq\det(\mathcal{E}^{\prime}_{i-1})\otimes\mathcal{O}_{X}(D^{\prime}_{i})$. By assumption (ii) there exists a hierarchical filtration of $\mathcal{E}^{\prime\prime}$ normalized by $\Lambda_{0}\otimes\det(\mathcal{E}^{\prime})$,
$\mathcal{E}^{\prime\prime}_{0}\subset\mathcal{E}^{\prime\prime}_{1}\subset\cdots\subset\mathcal{E}^{\prime\prime}_{b}=\mathcal{E}^{\prime\prime},$
with determinant increments $\mathcal{O}_{X}(D^{\prime\prime}_{j})$.
Hypothesis (iii) ensures that each divisor $D^{\prime\prime}_{j}$ yields an elementary transform of $\mathcal{E}$ lifting the elementary transform of $\mathcal{E}^{\prime\prime}$: the reason is that, under disjoint support from the torsion and the universal property of elementary transforms (cf. Huybrechts–Lehn, Prop. 1.1.8), there exist saturated inclusions
$\mathcal{E}^{(j-1)}\subset\mathcal{E}^{(j)}\subset\mathcal{E}^{(j-1)}(D^{\prime\prime}_{j})$
with torsion quotient supported on $D^{\prime\prime}_{j}$, and whose pushforward to $\mathcal{E}^{\prime\prime}$ reproduces the corresponding step of the $\mathcal{E}^{\prime\prime}$–filtration.
Therefore, after exhausting the $\mathcal{E}^{\prime}$–filtration inside $\mathcal{E}$ using (A), we may lift each step of the $\mathcal{E}^{\prime\prime}$–filtration to $\mathcal{E}$, producing a chain
$\mathcal{E}_{0}\subset\mathcal{E}_{1}\subset\cdots\subset\mathcal{E}_{a+b}=\mathcal{E}$
of saturated same–rank inclusions. This is a legitimate hierarchical filtration of $\mathcal{E}$ of length $a+b$. Thus
@@
$h_{\Lambda_{0}}(\mathcal{E})\geq a+b=h_{\Lambda_{0}}(\mathcal{E}^{\prime})+h_{\Lambda_{0}\otimes\det(\mathcal{E}^{\prime})}(\mathcal{E}^{\prime\prime}),$
and (3) follows immediately.
(D) If $\mathcal{E}=\mathcal{E}_{1}\oplus\mathcal{E}_{2}$ and both bundles admit maximal hierarchical filtrations whose determinant increments correspond to divisors with pairwise disjoint support, then elementary transforms with disjoint support commute, so the two filtrations can be interleaved to produce a filtration of $\mathcal{E}$ whose length is the sum of the two lengths. The claimed inequality follows. If supports intersect, such commutativity need not hold, and ranks may drop under saturation, so the simple additive formula may fail. ∎
@@ 5. Effects of the Minimal Model Program on hierarchical filtrations
In this section we study the behavior of hierarchical filtrations of vector bundles under birational transformations arising in the Minimal Model Program (MMP) when the base variety has dimension at most $2$. Since the MMP for smooth varieties of dimension $\leq 2$ preserves smoothness, we may restrict throughout to smooth projective curves and surfaces.
@@ 5.5. Hierarchical depth under MMP
Let $X_{\min}=\mathbb{P}^{2}$ with hyperplane class $H$, and let $f:X\to X_{\min}$ be the blowup of $\mathbb{P}^{2}$ at two distinct points $p_{1},p_{2}$. Denote by $E_{1},E_{2}\subset X$ the corresponding exceptional divisors. Then $\mathrm{Pic}(X)=\mathbb{Z}\langle f^{*}H\rangle\oplus\mathbb{Z}\langle E_{1}\rangle\oplus\mathbb{Z}\langle E_{2}\rangle$.
Fix the following data:
@@
Then $\det(\mathcal{E}_{\min})\cong\mathcal{O}_{\mathbb{P}^{2}}(5)$, and by Proposition 4.1 we have
@@
Since $\mathrm{rank}(\mathcal{E})=2$ we compute the determinant on $X$ as
@@
Take the normalization on $X$ to be $\Lambda_{0}=\mathcal{O}_{X}(E_{2})$; equivalently $\Lambda_{0}\simeq f^{*}\Lambda_{0,\min}\otimes\mathcal{O}_{X}(0\cdot E_{1}+1\cdot E_{2})$. Thus in the notation of Theorem 5.6 we have the integer coefficients
$\alpha_{1}=2,\ \alpha_{2}=4,\qquad\beta_{1}=0,\ \beta_{2}=1,$
and indeed $\alpha_{j}-\beta_{j}\geq 0$ for $j=1,2$.
We now compute both sides of formula (5). On the minimal model we already have $h_{\Lambda_{0,\min}}(\mathcal{E}_{\min})=5$. On $X$ the class
$\Delta=\det(\mathcal{E})\otimes\Lambda_{0}^{-1}\simeq f^{*}\mathcal{O}_{\mathbb{P}^{2}}(5)\otimes\mathcal{O}_{X}(2E_{1}+3E_{2})$
is represented by an effective divisor of class $5f^{*}H+2E_{1}+3E_{2}$. One may decompose a representative of this class as the sum of
  • five irreducible strict transforms of lines in $\mathbb{P}^{2}$ (each in the class $f^{*}H$), and
  • two copies of $E_{1}$ and three copies of $E_{2}$.
Thus the maximal possible number of nonzero irreducible effective summands in $\Delta$ equals
$5+2+3=10.$
By the definition of hierarchical depth (maximal number of effective divisorial increments for $\det(\mathcal{E})\otimes\Lambda_{0}^{-1}$) this shows $h_{\Lambda_{0}}(\mathcal{E})=10$.
Comparing with Theorem 5.6 we have
$h_{\Lambda_{0}}(\mathcal{E})=h_{\Lambda_{0,\min}}(\mathcal{E}_{\min})+\sum_{j=1}^{2}(\alpha_{j}-\beta_{j})=5+(2-0)+(4-1)=10,$
in perfect agreement with the direct decomposition above.
@@ 6.1. Slope behavior under hierarchical increments
and for each index $i$ the inequality $\mu_{H}(\mathcal{E}_{i})>\mu_{H}(\mathcal{E}_{i-1})$ holds precisely when the corresponding determinant increment $D_{i}$ is nonzero ($D_{i}\not\sim 0$).
@@ 6.3. Comparison with Harder–Narasimhan filtrations
Split bundle on a curve
Let $C$ be a smooth projective curve over a field and set
@@ 7. Applications to algebraic–geometric codes
In the study of constructing algebraic geometry (AG) codes on varieties, it is standard to assume that the underlying variety is smooth and irreducible, since the definition of such codes relies on the well-behaved structure of the function field and divisor theory on smooth varieties. However, within the framework of the Minimal Model Program (MMP), the situation becomes more delicate: the birational operations used in the MMP—such as divisorial contractions and flips—may introduce singularities, even if one starts with a smooth variety. In particular, in dimensions three and higher, the MMP involves both divisorial contractions (which contract subvarieties of codimension one) and flips (which replace certain singular loci by others), and these operations generally lead to varieties with terminal or canonical singularities. Consequently, to maintain smoothness throughout the process, one must restrict attention to varieties of dimension at most two. For smooth projective surfaces, the MMP simplifies considerably: there are no flips, and the only possible operations are contractions of (1)(-1)-curves. Each such contraction is the inverse of a blowup of a smooth point and therefore preserves smoothness. Hence, in the case of smooth surfaces, the MMP coincides with the classical Castelnuovo contraction process, ensuring that the minimal model obtained remains smooth. This restriction to smooth varieties (and thus to surfaces) is essential for the consistent application of AG code constructions on varieties governed by the MMP framework.
@@ 7.2. Minimal models and coding–theoretic optimization
Let $f:Y\to X$ be the contraction of a (1)(-1)–curve $E\subset Y$ onto a smooth $\mathbb{F}_{q}$–rational point $P_{0}\in X$. Let $\mathcal{E}_{Y}$ be a vector bundle on $Y$ whose hierarchical filtration has no effective increment supported on $E$, and set $\mathcal{E}:=f_{*}\mathcal{E}_{Y}$ (after torsion saturation). Then there is a canonical identification
@@
and for any evaluation set $D\subset X(\mathbb{F}_{q})\setminus\{P_{0}\}$ one has a commutative diagram
[…]
Consequently,
@@
Since $f$ is a birational morphism contracting a smooth rational curve with normal bundle $\mathcal{O}_{E}(-1)$, we have $R^{i}f_{*}\mathcal{O}_{Y}=0$ for $i>0$. If no increment of the hierarchical filtration of $\mathcal{E}_{Y}$ is supported on $E$, then $\mathcal{E}_{Y}$ is numerically trivial along $E$, and thus $R^{i}f_{*}\mathcal{E}_{Y}=0$ for $i>0$. By the projection formula,
$H^{0}(Y,\mathcal{E}_{Y})\cong H^{0}(X,f_{*}\mathcal{E}_{Y})=H^{0}(X,\mathcal{E}),$
which yields the claimed isomorphism of global sections. For any evaluation divisor $D\subset X(\mathbb{F}_{q})\setminus\{P_{0}\}$, the points $f^{-1}(D)$ correspond bijectively to $D$, and the fibers $\mathcal{E}|_{P_{i}}\cong\mathcal{E}_{Y}|_{f^{-1}(P_{i})}$ are canonically identified. Hence the evaluation maps coincide, giving the commutative diagram and equality of codes. ∎
@@
Let $X_{\min}$ be the minimal model of a smooth projective surface $X$ over $\mathbb{F}_{q}$, and let $f:X\to X_{\min}$ denote the composition of its (1)(-1)–curve contractions. For every vector bundle $\mathcal{E}$ on $X$ whose hierarchical filtration has no effective increment supported on the exceptional locus of $f$, one has
$C_{X}(\mathcal{E},D)=C_{X_{\min}}(f_{*}\mathcal{E},\,f(D)).$
In particular, all essential code parameters (length, dimension, and minimum distance) are preserved under the MMP reduction to the minimal model.
In the following we show that MMP can improve code parameters.
Proposition 7.3.
Let $f:Y\to X$ be the contraction of a set of (1)(-1)-curves $E_{1},\dots,E_{m}$ on a smooth projective surface $Y$, and let $\mathcal{E}_{Y}$ be a vector bundle of rank $r$ on $Y$ with $H^{0}(Y,\mathcal{E}_{Y})$ of dimension $k$. Set $\mathcal{E}:=f_{*}\mathcal{E}_{Y}$.
Let $D_{Y}\subset Y(\mathbb{F}_{q})$ be a set of $N_{Y}$ evaluation points, and let $D_{E}=D_{Y}\cap\bigcup_{i=1}^{m}E_{i}$ be the subset of points lying on the exceptional divisors, with $D_{E}=N_{E}$. Let $D=f(D_{Y}\setminus D_{E})\subset X$.
Define the AG codes:
$C_{Y}:=\operatorname{ev}_{D_{Y}}(H^{0}(Y,\mathcal{E}_{Y}))\subset(\mathbb{F}_{q}^{r})^{N_{Y}},\quad C_{X}:=\operatorname{ev}_{D}(H^{0}(X,\mathcal{E}))\subset(\mathbb{F}_{q}^{r})^{N_{Y}-N_{E}}.$
Then:
  1. (i)
    The dimension is preserved under contraction:
    $\dim C_{X}=\dim C_{Y}=k.$
  2. (ii)
    The absolute minimum distance is also preserved:
    $d_{\min}(C_{X})=d_{\min}(C_{Y})=:d_{\min}.$
  3. (iii)
    The normalized minimum distance improves according to
    $\delta_{Y}:=\frac{d_{\min}}{r\,N_{Y}}\leq\delta_{X}:=\frac{d_{\min}}{r\,(N_{Y}-N_{E})}.$
  4. (iv)
    In particular, the improvement factor of relative distance is
    $\frac{\delta_{X}}{\delta_{Y}}=\frac{N_{Y}}{N_{Y}-N_{E}}>1,$
    which grows with the proportion of evaluation points lying on exceptional divisors.

### v3

@@ 4. Hierarchical depth and determinant bounds
(A) Let $\mathcal{E}^{\prime}_{0}\subset\cdots\subset\mathcal{E}^{\prime}_{h^{\prime}}=\mathcal{E}^{\prime}$ be a hierarchical filtration of $\mathcal{E}^{\prime}$ normalized by $\Lambda_{0}$. Since $\mathcal{E}^{\prime}\subset\mathcal{E}$ is saturated of the same rank, each $\mathcal{E}^{\prime}_{i}\subset\mathcal{E}$ remains saturated and the quotients $\mathcal{E}/\mathcal{E}^{\prime}_{i}$ are torsion with the same support as $\mathcal{E}^{\prime}/\mathcal{E}^{\prime}_{i}$. Thus the same chain is a hierarchical filtration of $\mathcal{E}$, yielding $h_{\Lambda_{0}}(\mathcal{E})\geq h^{\prime}$. Taking suprema gives $h_{\Lambda_{0}}(\mathcal{E}^{\prime})\leq h_{\Lambda_{0}}(\mathcal{E})$.
[…]
Step 1: Embedding the filtration of $\mathcal{E}^{\prime}$.
By (A) (monotonicity), the hierarchical filtration $\mathfrak{F}^{\prime}$ of $\mathcal{E}^{\prime}$ embeds into $\mathcal{E}$: each $\mathcal{E}^{\prime}_{i}\subset\mathcal{E}^{\prime}\subset\mathcal{E}$ is saturated in $\mathcal{E}$ and the quotients $\mathcal{E}/\mathcal{E}^{\prime}_{i}$ are torsion with the same support as $\mathcal{E}^{\prime}/\mathcal{E}^{\prime}_{i}$. Thus we obtain a chain of saturated subbundles of $\mathcal{E}$:
$\mathcal{E}^{\prime}_{0}\subset\mathcal{E}^{\prime}_{1}\subset\cdots\subset\mathcal{E}^{\prime}_{a}=\mathcal{E}^{\prime}\subset\mathcal{E}.$
Each inclusion $\mathcal{E}^{\prime}_{i-1}\subset\mathcal{E}^{\prime}_{i}$ remains an elementary transform along some effective divisor $C_{i}$, and $\det(\mathcal{E}^{\prime}_{0})\simeq\Lambda_{0}$ by hypothesis.
Step 2: Lifting the filtration of $\mathcal{E}^{\prime\prime}$ step by step.
Set $\mathcal{F}_{0}:=\mathcal{E}^{\prime}$ (the last term of (2)). By assumption, for $j=1$ we have a diagram ($\star_{1}$) with $\mathcal{F}_{0}=\mathcal{E}^{\prime}$ and $\mathcal{F}_{1}\subset\mathcal{E}$ a saturated subbundle fitting into
$0\to\mathcal{E}^{\prime}\to\mathcal{F}_{1}\to\mathcal{E}^{\prime\prime}_{1}\to 0,$
and $\mathcal{F}_{0}\subset\mathcal{F}_{1}$ is an elementary transform along $D_{1}$. The determinant satisfies
$\det(\mathcal{F}_{1})\simeq\det(\mathcal{F}_{0})\otimes\mathcal{O}_{X}(D_{1})\simeq\det(\mathcal{E}^{\prime})\otimes\mathcal{O}_{X}(D_{1}).$
Proceed inductively: having constructed $\mathcal{F}_{j-1}\subset\mathcal{E}$ with $0\to\mathcal{E}^{\prime}\to\mathcal{F}_{j-1}\to\mathcal{E}^{\prime\prime}_{j-1}\to 0$, diagram ($\star_{j}$) provides $\mathcal{F}_{j}\subset\mathcal{E}$ containing $\mathcal{F}_{j-1}$ as an elementary transform along $D_{j}$, and $0\to\mathcal{E}^{\prime}\to\mathcal{F}_{j}\to\mathcal{E}^{\prime\prime}_{j}\to 0$. The determinant relation is
$\det(\mathcal{F}_{j})\simeq\det(\mathcal{F}_{j-1})\otimes\mathcal{O}_{X}(D_{j}).$
Step 3: Concatenation yields a hierarchical filtration of $\mathcal{E}$.
Chaining (2) with the lifts of the $\mathcal{E}^{\prime\prime}$-filtration gives a sequence
$\mathcal{E}^{\prime}_{0}\subset\cdots\subset\mathcal{E}^{\prime}_{a}=\mathcal{F}_{0}\subset\mathcal{F}_{1}\subset\cdots\subset\mathcal{F}_{b}=:\widetilde{\mathcal{E}}$.
We claim $\widetilde{\mathcal{E}}$$\mathcal{E}$. Indeed, $\mathcal{F}_{b}$ fits into $0\to\mathcal{E}^{\prime}\to\mathcal{F}_{b}\to\mathcal{E}^{\prime\prime}_{b}=\mathcal{E}^{\prime\prime}\to 0$, and the inclusion $\mathcal{F}_{b}\subset\mathcal{E}$ induces the identity on $\mathcal{E}^{\prime}$ and the given surjection $\mathcal{E}\to\mathcal{E}^{\prime\prime}$. Hence $\mathcal{F}_{b}=\mathcal{E}$ by the five‑lemma (or by comparing ranks and the fact that both are subbundles of $\mathcal{E}$).
Each inclusion in (4) is saturated of rank $r$ and the successive quotients are torsion sheaves supported on effective divisors ($C_{i}$ for the $\mathcal{E}^{\prime}$-part, $D_{j}$ for the $\mathcal{E}^{\prime\prime}$-part). Moreover, from (3) and the normalization of $\mathfrak{F}^{\prime\prime}$ we have
@@
$\det(\mathcal{F}_{j})\simeq\det(\mathcal{E}^{\prime})\otimes\det(\mathcal{E}^{\prime\prime}_{j})\simeq\det(\mathcal{E}^{\prime})\otimes\bigl(\Lambda_{0}^{\prime}\otimes\mathcal{O}_{X}(D_{1}+\cdots+D_{j})\bigr)\simeq\Lambda_{0}\otimes\mathcal{O}_{X}(D_{1}+\cdots+D_{j}),$
so the determinant increments match those of $\mathfrak{F}^{\prime\prime}$. Consequently, (4) is a hierarchical filtration of $\mathcal{E}$ normalized by $\Lambda_{0}$, of total length $a+b$.
Step 4: Conclusion.
Since $a=h_{\Lambda_{0}}(\mathcal{E}^{\prime})$ and $b=h_{\Lambda_{0}^{\prime}}(\mathcal{E}^{\prime\prime})$ by definition of hierarchical depth, the existence of the filtration (4) shows
$h_{\Lambda_{0}}(\mathcal{E})\geq a+b=h_{\Lambda_{0}}(\mathcal{E}^{\prime})+h_{\Lambda^{\prime}_{0}}(\mathcal{E}^{\prime\prime}),$
which is precisely inequality (1).
(C) Under the disjoint support hypothesis, Lemma 3.2 ensures the elementary transforms from the two filtrations commute. Interleaving them produces a hierarchical filtration of $\mathcal{E}$ of length $h_{\Lambda_{0}}(\mathcal{E}_{1})+h_{\Lambda_{0}}(\mathcal{E}_{2})$, giving equality. If supports intersect, commutativity may fail and saturation can reduce the achievable length, yielding only the inequality. ∎
Remark 4.4.
(Non-monotonicity for smaller rank) If $\mathcal{E}^{\prime}\subset\mathcal{E}$ is a subbundle with $\mathrm{rank}(\mathcal{E}^{\prime})<\mathrm{rank}(\mathcal{E})$, the inequality $h_{\Lambda_{0}}(\mathcal{E}^{\prime})\leq h_{\Lambda_{0}}(\mathcal{E})$ may fail. For example, on $X=\mathbb{P}^{1}$ with $\Lambda_{0}=\mathcal{O}_{\mathbb{P}^{1}}$, take $\mathcal{E}^{\prime}=\mathcal{O}(1)\subset\mathcal{E}=\mathcal{O}(1)\oplus\mathcal{O}(-1)$. Then $h(\mathcal{E}^{\prime})=\deg\mathcal{O}(1)=1$ while $h(\mathcal{E})=\deg\det(\mathcal{E})=0$.
@@ 5. Effects of the Minimal Model Program on hierarchical filtrations
In this section we study the behavior of hierarchical filtrations of vector bundles under birational transformations arising in the Minimal Model Program (MMP) when the base variety has dimension at most $2$. For smooth projective surfaces, the MMP consists solely of contractions of (1)(-1)-curves, which preserve smoothness.
@@ 5.5. Hierarchical depth under MMP
Let $X_{\min}=\mathbb{P}^{2}$ with hyperplane class $H$, and let $f:X\to X_{\min}$ be the blowup of $\mathbb{P}^{2}$ at two distinct points $p_{1},p_{2}$. Denote by $E_{1},E_{2}\subset X$ the corresponding exceptional divisors. Then
$\mathrm{Pic}(X)=\mathbb{Z}\langle f^{*}H\rangle\oplus\mathbb{Z}\langle E_{1}\rangle\oplus\mathbb{Z}\langle E_{2}\rangle.$
Fix the following data on $X_{\min}$:
@@
Then $\det(\mathcal{E}_{\min})\cong\mathcal{O}_{\mathbb{P}^{2}}(5)$. For a split bundle on $\mathbb{P}^{2}$, Proposition 2.4 (with $A=H$, $d=5$, $d_{0}=0$) gives the bound
$h_{\Lambda_{0,\min}}(\mathcal{E}_{\min})\leq 5.$
In fact, since $\mathbb{P}^{2}$ has Picard rank one and the class $H$ admits pairwise disjoint representatives (distinct lines), one can achieve equality:
@@
Since $\mathrm{rank}(\mathcal{E})=2$, we compute
@@
Take the normalization on $X$ to be
$\Lambda_{0}=\mathcal{O}_{X}(E_{2})\simeq f^{*}\mathcal{O}_{\mathbb{P}^{2}}\otimes\mathcal{O}_{X}(E_{2}).$
Thus in the notation of Theorem 5.6 we have:
$\det(\mathcal{E})\simeq f^{*}\det(\mathcal{E}_{\min})\otimes\mathcal{O}_{X}(\alpha_{1}E_{1}+\alpha_{2}E_{2}),\quad\alpha_{1}=2,\ \alpha_{2}=4,$
$\Lambda_{0}\simeq f^{*}\Lambda_{0,\min}\otimes\mathcal{O}_{X}(\beta_{1}E_{1}+\beta_{2}E_{2}),\quad\beta_{1}=0,\ \beta_{2}=1.$
Now Theorem 5.6 predicts:
$h_{\Lambda_{0}}(\mathcal{E})=h_{\Lambda_{0,\min}}(\mathcal{E}_{\min})+(\alpha_{1}-\beta_{1})+(\alpha_{2}-\beta_{2})=5+(2-0)+(4-1)=10.$
To verify directly, consider the class
$\Delta:=\det(\mathcal{E})\otimes\Lambda_{0}^{-1}\simeq f^{*}\mathcal{O}_{\mathbb{P}^{2}}(5)\otimes\mathcal{O}_{X}(2E_{1}+3E_{2}).$
A representative of $\Delta$ can be decomposed as:
$\Delta\sim(f^{*}H)+\cdots+(f^{*}H)+E_{1}+E_{1}+E_{2}+E_{2}+E_{2},$
where we have 5 copies of $f^{*}H$ (strict transforms of lines not passing through $p_{1}$ or $p_{2}$), 2 copies of $E_{1}$, and 3 copies of $E_{2}$. Each summand gives a possible determinant increment in a hierarchical filtration, so the maximal length is $5+2+3=10$. Hence
$h_{\Lambda_{0}}(\mathcal{E})=10,$
confirming the theorem.
@@ 6.1. Slope behavior under hierarchical increments
and for each index $i$ the inequality $\mu_{H}(\mathcal{E}_{i})>\mu_{H}(\mathcal{E}_{i-1})$ holds precisely when $D_{i}$ is not numerically equivalent to zero. ($D_{i}\not\sim 0$).
@@ 6.3. Comparison with Harder–Narasimhan filtrations
Split bundle on a curve. Let $C$ be a smooth projective curve over a field and set
@@ 7. Applications to algebraic–geometric codes
Let $X$ be a smooth projective variety over a finite field $\mathbb{F}_{q}$, $\mathcal{E}$ a vector bundle on $X$, and $D\subset X(\mathbb{F}_{q})$ a set of $N$ rational points. The associated algebraic–geometric code is $C_{X}(\mathcal{E},D)=\operatorname{ev}_{D}(H^{0}(X,\mathcal{E}))\subset\mathbb{F}_{q}^{\,rN}$ where $r=\mathrm{rank}(\mathcal{E})$.
Hierarchical filtrations of $\mathcal{E}$ expose degeneracies: if an elementary transform occurs along a divisor containing evaluation points, the corresponding code coordinates vanish. Consequently, hierarchical depth $h_{\Lambda_{0}}(\mathcal{E})$ bounds the number of such degenerate layers.
For surfaces, the minimal model program contracts (1)(-1)-curves—smooth point blowups in reverse—and Theorem 5.6 quantifies how hierarchical depth changes under these birational maps. Contracting exceptional divisors eliminates the associated degenerate coordinates, improving the code’s normalized minimum distance while preserving dimension. The next subsections provide explicit constructions demonstrating this optimization.
@@ 7.2. Minimal models and coding–theoretic optimization
Let $f:Y\to X$ be the contraction of a (1)(-1)-curve $E\subset Y$ onto a smooth $\mathbb{F}_{q}$-rational point $P_{0}\in X$. Let $\mathcal{E}_{Y}$ be a vector bundle on $Y$ such that:
  1. (i)
    $\mathcal{E}_{Y}$ is $f$-trivial in the sense that $R^{1}f_{*}\mathcal{E}_{Y}=0$;
  2. (ii)
    $\det(\mathcal{E}_{Y})$ has no component supported on $E$ (equivalently, in the decomposition $\det(\mathcal{E}_{Y})\simeq f^{*}L\otimes\mathcal{O}_{Y}(mE)$, we have $m=0$).
Set $\mathcal{E}:=(f_{*}\mathcal{E}_{Y})^{\mathrm{sat}}$. Then there is a canonical identification
@@
and for any evaluation set $D\subset X(\mathbb{F}_{q})\setminus\{P_{0}\}$ one has
@@
Condition (i) gives $R^{i}f_{*}\mathcal{E}_{Y}=0$ for $i>0$ (since $R^{i}f_{*}\mathcal{E}_{Y}$ are supported at $P_{0}$ and vanish for $i>1$ by dimension reasons). The projection formula and the fact that $f_{*}\mathcal{O}_{Y}\cong\mathcal{O}_{X}$ yield
$H^{0}(Y,\mathcal{E}_{Y})\cong H^{0}(X,f_{*}\mathcal{E}_{Y})\cong H^{0}(X,\mathcal{E}),$
where the second isomorphism uses that saturation does not change global sections.
Condition (ii) ensures that for any $P\in D$, the natural map $(f_{*}\mathcal{E}_{Y})_{P}\to\mathcal{E}_{Y}|_{f^{-1}(P)}$ is an isomorphism (the torsion of $f_{*}\mathcal{E}_{Y}$ is concentrated at $P_{0}$). Hence evaluation at $P$ and at $f^{-1}(P)$ agree, giving the equality of codes. ∎
Remark 7.2.
Condition (i) holds, for instance, when $\mathcal{E}_{Y}|_{E}$ is a direct sum of line bundles of degree $\geq-1$, or when $\mathcal{E}_{Y}$ is pulled back from $X$ twisted by $\mathcal{O}_{Y}(-mE)$ with $0\leq m\leq 1$. Condition (ii) is exactly the requirement that no step in a hierarchical filtration of $\mathcal{E}_{Y}$ involves an elementary transform along $E$.
Remark 7.3.
While it is folklore that birational varieties can yield equivalent AG codes when evaluation points avoid exceptional loci, Proposition 7.1 provides precise cohomological and determinantal criteria in terms of hierarchical filtrations. In particular, condition (ii) is equivalent to the vanishing of the exceptional contribution $\delta$ in Theorem 5.6.
@@
Let $X$ be a smooth projective surface over $\mathbb{F}_{q}$, $f:X\to X_{\min}$ the contraction of all (1)(-1)-curves to its minimal model, and $\mathcal{E}$ a vector bundle on $X$ satisfying the hypotheses of Proposition 7.1 for each contraction step. Let $D\subset X(\mathbb{F}_{q})$ be an evaluation set, and decompose $D=D_{\mathrm{ex}}\cup D_{\mathrm{reg}}$ where $D_{\mathrm{ex}}$ lies on exceptional curves and $D_{\mathrm{reg}}$ does not. Set $D_{\min}:=f(D_{\mathrm{reg}})\subset X_{\min}(\mathbb{F}_{q})$.
Then the AG codes are related by
$C_{X}(\mathcal{E},D)\cong C_{X_{\min}}(f_{*}\mathcal{E},\,D_{\min})\times\{0\}^{r\cdot|D_{\mathrm{ex}}|},$
where $\cong$ denotes isomorphism of $\mathbb{F}_{q}$-vector spaces. Consequently:
  • $\dim C_{X}(\mathcal{E},D)=\dim C_{X_{\min}}(f_{*}\mathcal{E},D_{\min})$;
  • $d_{\min}(C_{X}(\mathcal{E},D))=d_{\min}(C_{X_{\min}}(f_{*}\mathcal{E},D_{\min}))$;
  • $|D|=|D_{\min}|+|D_{\mathrm{ex}}|$, so the length decreases by $r\cdot|D_{\mathrm{ex}}|$ coordinates.
In particular, the normalized minimum distance $\delta=d_{\min}/(r\cdot|D|)$ strictly increases when $D_{\mathrm{ex}}\neq\varnothing$.
Corollary 7.5.
Under the hypotheses of Proposition 7.1, let $D_{\mathrm{ex}}\subset D$ be evaluation points on exceptional curves. Then
$\frac{d_{\min}(C_{X_{\min}}(f_{*}\mathcal{E},f(D\setminus D_{\mathrm{ex}})))}{r\cdot|D\setminus D_{\mathrm{ex}}|}>\frac{d_{\min}(C_{X}(\mathcal{E},D))}{r\cdot|D|},$
with strict inequality when $D_{\mathrm{ex}}\neq\varnothing$.